A flywheel is driven by a chain that travels 4ft./second. Find the diameter (in inches) of the wheel if it makes 144 rpm.
Answer by Theo(13342)

i think it works like this.
the chain travels at 4 feet per second.
multiply that by 12 and the chain is traveling at 48 inches per second.
multiply that by 60 and the chain is traveling at 2880 inches per minute.
in that minute, the flywheel makes 144 rpm.
take 2880 and divide it by 144 to get a circumference of 20 for the flywheel.
since circumference = 2 * pi * r and since diameter = 2 * r, then the circumference can be modeled as circumference = pi * diameter.
since circumference = 20, then solve for diameter to get diameter = 20 / pi = 6.366197724 inches.
that should be your answer, if i did this correctly.

you have circumference of 20.
in 144 revolutions of the flywheel, it has traveled 20 * 144 = 2880 inches.
since the chain is driving the flywheel, then the circumference of the flywheel should be traveling the same distance as the chain.
it appears that this solution allows that to happen.